OpenStudy (mackenzie2013):

What is the value of a3 - b2 for a = 3 and b = 1/2 ?

OpenStudy (anonymous):

substitute it

OpenStudy (whpalmer4):

a^3 - b^2 = a*a*a - b*b plug in the values for a and b, what do you get?

OpenStudy (mackenzie2013):

9-1? so the answer would be 8. Right?

OpenStudy (whpalmer4):

1/2 * 1/2 = 1? are you sure?

OpenStudy (whpalmer4):

also, it's a*a*a, so 3*3*3, not 3*3

OpenStudy (anonymous):

is there no exponent? or just a whole num?

OpenStudy (mackenzie2013):

The answer is 26 3/4

OpenStudy (anonymous):

a3 - b2 if it just a whole number then the constant num should be first like 3a-2b hahah !! i'm just confused

OpenStudy (whpalmer4):

right! good job, @Mackenzie2013

OpenStudy (whpalmer4):

but in the future, to avoid confusion, you should write a^3, not a3

OpenStudy (anonymous):

hahha!! its exponent

OpenStudy (whpalmer4):

@pakkibu you misunderstood the problem. it is the equivalent of a*a*a - b*b, NOT 3a-2b

OpenStudy (shaik0124):

27-1/4 108-1/4 107/4 is answer